# Limits to Predicting Online Speech Using Large Language Models

### v1

@@ 4 Results
We divide the presentation of our main findings into three parts, one about the unpredictability of posts in general (Section 4.1), one about predicting from different sources of context (Section 4.2), and one about our investigation into the sources of unpredictability (Section 4.3).
Figure 4: Distribution of the average NLL of our subjects (with no additional context). As comparison we include the average NLL of three financial news accounts: YahooFinance, CNBC and Stocktwits. Colloquially financial news are hard to predict, therefore most of our subjects are too. Model: GPT-2-XL.
4.1 Unpredictability of Posts
As described in Section 3.3, we used negative log-likelihood as a measure of the model’s uncertainty when predicting a user’s tweets. In Figure 4 we present the distribution of the average NLL of our subjects. To gain an intuition on how hard it is to predict our subjects’ tweets relative to other accounts, we included three popular financial news accounts as comparison: YahooFinance, CNBC, and Stocktwits. Intuitively, predicting financial news and the stock market is hard Johnson et al. (2003), which should make them less predictable. However, Figure 4 reveals that most subjects in our pool are actually harder to predict than those news accounts. Of course, while the content of those tweets may be hard to predict, their style and vocabulary may not.
Finally, the absolute information content of users’ tweets (without additional context) is between 1.5-2 bits per character, depending on which model we use (see ’no context’ bars in Figure 2(c)). As a comparison, Shannon’s upper bound on the cross-entropy of the English language is 1.3 bits per character Shannon (1951), and recent estimates using neural language models say it is as low as 1.12 bits Takahashi and Tanaka-Ishii (2018). Only our largest model with additional user context achieves comparable entropy, with an average of 1.1193 bits per character (Llama-2 with user context). These results suggest that individual pieces of our online expression taken out of context are far from predictable using today’s LLMs.
@@ 4.1 Unpredictability of Posts
(b) Finetuning on combinations of contexts. For example ["peer", "random"] denotes that half of the tweets were sampled from $\mathcal{T}^{\texttt{peer}}_{u}$ and the other half from $\mathcal{T}^{\texttt{random}}_{u}$. Mixing contexts does not always lead to better results, suggesting overlap in predictive information.
@@
Figure 6: Average loss curves with standard error for finetuning experiments on GPT-2-XL. For each subject ($N=5102$) we finetune on the specified context, and compute the loss (negative log-likelihood) on $\mathcal{T}_{u}^{\texttt{eval}}$. The plotted averages are computed over the loss curves of 1000 subjects.
Figure 7: Effect size matrix on Llama-2. We plot the average effect size of $c_{2}$ relative to $c_{1}$ on user predictability. This is done by looking at the difference in predictability for each user $\Delta^{c1}_{c2}(u)=\bar{L}^{c1}_{u}-\bar{L}^{c2}_{u}$, where $\bar{L}^{c}_{u}$ is the average negative log-likelihood of user $u$ under context $c$. Plotted values are standardized mean differences (SMD) of $\Delta^{c1}_{c2}$. Darker green means greater improvement in prediction.
4.2 Predicting Posts from Context
We quantified how additional context influenced predictability. We used two methods of feeding additional information to our model, with complementary strengths and weaknesses: one prompting based approach and one finetuning based approach.
We start with the results of the prompting approach. In Figure 2, we plot the average BPC calculated over all of our subjects for three different models with varying parameter sizes. The no context case was consistently the most unpredictable with highest BPC. Depending on what context was included, the amount of improvement varied. We show that most of the predictive information is found in the user context, followed by peer and random context. This trend is consistent across all three of our models.
As part of our prompting experiments, we were also interested in quantifying the effect size each context had on predictability. We plot this for each context pair in Figure 7 on Llama-2. This effect size matrix illustrates our finding that all types of context have a large effect size ($>2.5\sigma$) over having no context at all (last column). It also allows us to quantify the effect size of our main finding: user context offers $1.8\sigma$ improvement over peer context — an even bigger improvement than what peer context offers over random context ($1.3\sigma$). While these effect sizes are most pronounced on Llama-2, we found that these results also translate to the other models as well (Figure 19 in Appendix).
@@ 4.2 Predicting Posts from Context
Our main finding was replicated in our finetuning experiments as well. Figure 6(a) shows the average loss curves (computed over 1000 subjects) for finetuning GPT-2-XL on different types of contexts. Again, for all three contexts the loss goes down significantly. However, the final loss they converge to is different, with large gaps between each. If we order contexts based on the achieved loss in the final round, we get the same order as before: user context is best, followed by peer context, then by the random control context. Bagrow et al. (2019)’s findings suggest that with enough peers, one can match the predictive information inside the user context (using 8-9 peers). However even with 15 peers per subject, we find no evidence of such peer-based user profiling.
@@
We established that user context always outperfors peer context, regardless of model choice or experimental method. Still, one might argue that by combining user and peer context, one might achieve even higher predictability than by using user context alone. This could suggest that there is some non-overlapping predictive information inside the peer context wrt. the user context. We test this hypothesis by finetuning on a mixture of contexts (Fig. 6(b)) to give us an idea which contexts shared common information. Combining peer+random contexts resulted in a linear interpolation of the final losses of finetuning on either context. In other words, mixing random and peer context outperformed the final loss of finetuning exclusively on random context by a large margin. Here, we found evidence of additional predictive information in the peer context, which was not contained in the random context. However, mixing peer+user contexts did not result in a significantly lower final loss. This suggests that there is a significant overlap in predictive information between peer and user context.
A final, clarifying note on the "jumps" in the loss curves: Since each $\mathcal{T}_{u}^{c}$ contains 250 tweets of varying lengths, 5 epochs of training resulted in different global steps for each user-context combination. This explains the irregularities around global steps 15, 20, … etc. since for some users we only have losses up to that global step.
Figure 8: Average improvement in NLL from additional user context (compared to none). The first few tokens of a tweet benefit most from the additional context. Model: Llama-2. We plot up to 100. token, because the vast majority of tweets have $<100$ tokens (Fig 15 in Appendix).
4.3 What Drives Unpredictability
Finally, we were curious where and how context improved predictability. Perhaps a mildly surprising observation we can make based on Figure 2 is that random context also improves predictability in a non-negligible way. Due to the randomized nature of the control group there should be very little connection to our subjects’ tweets semantically. However the improvement we observed can be explained by added information regarding the syntax of a tweet – how long it is, the presence of @-mentions and hashtags, etc. Indeed this turns out to be the case: with random context, the predictability of the @ token improved the most compared all other tokens – see Appendix Table 1.
Even more surprisingly, predicting @-mentions and hashtags correctly played a significant role even in the case of user and peer context. Visualizing the improvement in predictability over individual tokens inside a tweet we noticed an interesting phenomenon (Figure 8). The initial“spike” in the figure suggests that some of the most improved tokens were at the start of tweets, often tokens part of @-mentions. After that, the average improvement stabilizes, then slightly goes up towards the end of the tweets, where hashtags would often be located. Based on these results it seemed plausible that @-mentions and hashtags were locations of greatest uncertainty, where additional context helped model prediction the most.
We repeated our prompting experiments after removing @-mentions and hashtags from our tweets. Assuming these were indeed one of the greatest sources of error, removing them would make our tweets more predictable overall, and the effect of additional context on predictability less strong. The results of our repeated experiments can be seen in Figure 9. Indeed, subjects became more predictable on average, and the effect of context on predictability decreased (by $\sim 20\%$ for random context, $\sim 4\%$ for peer context and $\sim 7\%$ for user context on Llama-2 70b). One peculiar finding was that for Falcon additional context did not improve predictability the same way as before. For example, the effect size of the user context dropped to only $0.4\sigma$, which is significantly less than the $2.1\sigma$ from before. A possible explanation is that the only (useful) predictive signal Falcon was able to pick up on was in the removed pieces of text. Another possibility is that it is simply more sensitive to discontinuities inside the text than other models.
Figure 9: Rerunning our experiments without @-mentions and hashtags results in these changed predictabilities. Subjects become more predictable on average, and the positive effect of context on predictability decreases. For Falcon it almost completely vanishes (with only user context offering slight improvement). Lighter bars are the results from our original experiment for comparison.
@@ 5 Limitations
While negative log-likelihood is the most popular measure of model uncertainty, there are certain limitations to using this metric. One is that it is calculated exclusively on tokens contained in tweet $T$, and does not take into account any improvements outside of that token (i.e. other tokens that are close to it in the embedding space). Another issue is the sensitivity of this metric: to the prediction of the first token (see Fig. 8 and Appendix B.5) and to slight changes in prompting (how tweets were separated, for example). Using low-frequency tokens as a separator between tweets (such as the eos token, which is usually reserved for separating training documents) produced abnormally high NLLs, which is why we decided to use more common tokens, such as space or newline for the prompting experiments. Another important aspect is the length of input; NLL is commonly lower for longer pieces of text. This is true here as well, and we show the relationship between average tweet length and model uncertainty in Appendix B.2. While these points affect absolute numbers in our analysis, it does not affect our main statements about the relative comparisons between different models and contexts. As an example, our choice of model also had a considerable effect on the absolute value of BPC (with GPT-2-XL scoring typically higher than Llama-2), however it did not affect the relative relationship between the different settings.
@@
Furthermore, we cannot rule out the possibility that some tweets in our corpus were part of the training dataset of the LLMs that we used. However, we believe the risk and severity of data contamination to be limited. GPT-2 was trained on WebText with content up to December 2017 Radford et al. (2019) ($95\%$ of our subjects’ tweets were authored after 2022), Falcon was trained on RefinedWeb which is is built using all CommonCrawl dumps until the 2023-06 one Penedo et al. (2023) (CommonCrawl typically does not contain snapshots of Twitter com (2024)) and the authors of Llama-2 claim to have “made an effort to remove data from certain sites known to contain a high volume of personal information about private individuals.” Touvron et al. (2023).
@@ 6 Discussion
We presented the results of an investigation using state-of-the-art large language models into the predictability of online speech by analyzing posts on X (Twitter). As the basis of our study, we collected posts from more than five thousand users’ timelines and their peers. We used a total of 6.25M tweets for our prompting experiments, plus an additional 10M for finetuning.
We found that it is surprisingly difficult to predict online speech. The majority of our users’ tweets are less predictable than tweets about financial news. Even with additional context, prediction performs relatively poorly. Only our largest model with additional user context is able to approximate the estimated information content of the English language. Further analysis on what drives this unpredictability reveals that the source of greatest uncertainty are hashtags and @-mentions. In other words, what improves prediction the most with additional context are basic signals: specifically, guessing the correct hashtags and @-mentions. All in all, our findings suggest that despite the impressive capabilities of large language models in other areas, state-of-the-art models predict online speech rather poorly.
@@
Our findings may also inform work on privacy and other ethical issues on social networks. For example, threats such as shadow profiling or impersonation on a global scale may not be as acute as some feared with state-of-the-art language models. However, these risks may still exist on a case-by-case basis (we demonstrate individual variability in Figure 14 in the Appendix). Moreover, there may be other conceivable harms that don’t map cleanly to questions of predictability.
7 Future Work
We presented results on the predictability of english speaking users’ online speech. This is only the first step towards contextualizing what it means for individuals to live in a world of transformers-based language models. It still remains to be answered how these results might translate to different sub-communities. Extending our analysis to different sub-communities and languages would be an interesting avenue for future work. Similarly, we welcome additional work into the external validity of our results. It is unclear to what extent our findings would translate to other social media platforms.
Finally, we find that additional research into the downstream effects of predictable speech on privacy is necessary. One well known issue about LLMs is that they are known to memorize personally identifiable information (PII) Lukas et al. (2023). We are happy to see a shift in recent literature acknowledging that preventing data leakage is not equivalent to meaningful privacy guarantees in the natural language context Brown et al. (2022).
@@ A.1.1 Sampling stage
This collection phase resulted in a pool of 10M tweets, with 5M unique authors. We sampled our subjects from this pool of authors ($N=5102$ sample). Strictly speaking, our sample will be biased towards users that have 1) been more active during our initial collection period and 2) are more active users in general. This was the closest we could get to a random sample with the offered API endpoints. It is also follows Bagrow et al. (2019)’s method of randomly sampling users. Again, following in their steps we dropped subjects that had scored high using the Bot-O-Meter API. Bot-O-Meter uses classifies users on a scale from 0-1 based on 200 of their tweets. See Figure 10 for a distribution of these scores. We dropped users that had a score higher than 0.5. Additionally, we dropped users that had a high retweet ratio (more than $80\%$ of their tweets consisted of retweets). This is an additional measure to prevent bot accounts in our subject pool (bots are known for frequently retweeting content Yang et al. (2020); Gilani et al. (2017)) as well as a practical consideration since we only wanted to include non-retweets in our analysis.
@@ A.1.2 Timeline collection stage
We used the Twitter API’s Timeline endpoint to query the subjects’ most recent tweets. To get tweets from around the same timeframe, we choose an end_time (which was the start of our sampling stage, 20. January). Tweets made after end_time were not included. Per subject, we collected a total of 500 tweets, half of which was reserved for evaluation, the other half served as user context. From these tweets, we identified the top-15 most mentioned users (the user’s peers) and collected 50 tweets from their timelines as well. These, as well as the user context tweets were collected such that they were authored before the oldest evaluation tweet ($t^{*}$ in Figure 12). From these, we selected the 250 most recent tweets. A simpler approach would have been to select exactly $\frac{250}{15}\approx 17$ tweets per peer. However, not all peers had this many tweets on their timeline before $t^{*}$, hence our choice for collecting more tweets per peer than necessary.

### v2

@@ 4 Results
Figure 4: Average effect size of context $c_{2}$ relative to $c_{1}$ on user predictability. Darker green means greater improvement in prediction. Model: Llama-2-70B.
In our experiments, we quantify how additional context influences model uncertainty, i.e. predictability. We used two methods of feeding additional information to our model, with complementary strengths and weaknesses: one prompting based approach which evaluates in-context learning and one finetuning based approach. We start with evaluating how in-context learning performs on base models (Section 4.1), and what they learn from the additional context (Section 4.2). Next, we evaluate tweet-tuned models, where we replicate our main findings by finetuning on different types of contexts (Section 4.3). Finally, we conclude that it is surprisingly hard to predict user tweets using large language models, even with additional context (Section 4.4).
4.1 Added Context Reduces Uncertainty in Base Models
We start with the results of the prompting approach on base models. In Figure 2, we plot the average BPC calculated over all of our subjects for three different models with varying parameter sizes. The no context case was consistently the most unpredictable with highest BPC. Depending on what context was included, the amount of improvement varied. We show that most of the predictive information is found in the user context, followed by peer and random context. This trend is consistent across all three of our models. For users that are hard to predict, we observe slightly larger improvements in model uncertainty (App. B.4).
We also quantify the effect size each context had on predictability. We plot this for each context pair in Figure 4 on Llama-2. This effect size matrix illustrates our finding that all types of context have a large effect size ($>2.5\sigma$) over having no context at all (last column). It also allows us to quantify the effect size of our main finding: user context offers $1.8\sigma$ improvement over peer context — an even bigger improvement than what peer context offers over random context ($1.3\sigma$). While these effect sizes are most pronounced on Llama-2, we found that these results also translate to our other models (Figure 22 in Appendix).
Figure 5: Average improvement in NLL from additional user context (compared to none). The first few tokens of a tweet benefit most from the additional context. Model: Llama-2-70B.
Figure 6: Average model uncertainty on tweets without @-mentions and hashtags. Subjects become more predictable on average, and the positive effect of context on predictability decreases. Lighter bars are the results from our original experiment for comparison.
Figure 7: Average model uncertainty on base vs. tweet-tuned models. First token was left out of the analysis (see Appendix B.6 for more detail). After tweet-tuning, in-context learning on additional context offers little to no improvement.
4.2 Base Models Learn Syntax From Added Context
Next, we were curious where and how context improves predictability. Interestingly, even random context improves predictability in a non-negligible way. We found evidence of the model learning to use the right syntax; i.e. how long tweets are, the presence of @-mentions and hashtags, etc. Indeed, the predictability of the ’@’ token improved the most compared to all other tokens (App. B.7) when using random context.
Predicting @-mentions and hashtags correctly plays a significant role even in the case of user and peer context. Visualizing the improvement in predictability over individual tokens inside a tweet we noticed an interesting phenomenon (Figure 5). Locations of greatest improvement were typically at the start and at the end of tweets – where @-mentions and hashtags would often be located. Assuming these were indeed one of the greatest sources of error in the no context case, removing them would make our tweets more predictable overall, and the effect of additional context on predictability less pronounced. Indeed, in Figure 6 subjects became more predictable on average, and the effect of context on predictability decreased (by $\sim 20\%$ for random context, $\sim 4\%$ for peer context and $\sim 7\%$ for user context on Llama-2 70b). In other words, models learned to assign higher probability to frequently used @-mentions and hashtags from context. While in some cases context ceases to be useful, the relative comparisons between different types of contexts remain the same.
@@ 4.2 Base Models Learn Syntax From Added Context
(b) Finetuning on mixtures of contexts. Mixing contexts does not always lead to better results, suggesting overlap in predictive information.
@@
Figure 8: Average loss curves with standard error for finetuning experiments on GPT-2-XL-tt. For each subject ($N=5102$) we finetune on the specified context, and compute the cross-entropy loss on $\mathcal{T}_{u}^{\texttt{eval}}$. The plotted averages are computed over the loss curves of 1000 subjects.
4.3 Results on Tweet-Tuned Models
Next, we prompt on tweet-tuned models. We plot the model uncertainty before and after tweet tuning for GPT-2-XL and Llama-3-8B in Figure 7. For both, tweet-tuning lowers model uncertainty in the no context case. However, predictability does not improve anymore with additional context in most cases. This is in line with our observations from before: base models learn typical twitter syntax from context. Our tweet-tuned models learn this through finetuning, limiting the amount of useful information that can be leveraged from in-context learning. For GPT-2-XL-tt, we find limited to no improvement through additional context; only Llama-3-8B-tt is able to learn more from additional user context. To overcome these limitations we observed in in-context learning, we instead turn to finetuning tweet tuned models on the context.
@@ 4.3 Results on Tweet-Tuned Models
Figure 8(a) shows the average loss curves (computed over 1000 subjects) for finetuning GPT-2-XL-tt on different types of contexts  [^1] A clarifying note on the ”jumps” in the loss curves (at steps 15, 20, … etc.): Since each $\mathcal{T}_{u}^{c}$ contains 250 tweets of varying lengths, 5 epochs of training resulted in different global steps for each user-context combination.. Again, for all three contexts the loss goes down significantly. However, the final loss they converge to is different, with large gaps between each. If we order contexts based on the achieved loss in the final round, we get the same order as before: user context is best, followed by peer context, then random context. Combined with the observed stability of user rankings across models (App. B.1), we believe this result would extend to larger models as well.
@@
We established that user context always outperforms peer context, regardless of model choice or experimental method. Still, one might argue that by combining user and peer context, one might achieve better results. This would suggest that there is some non-overlapping predictive information inside the peer context wrt. the user context. We test this hypothesis by finetuning on a mixture of contexts (Fig. 8(b)). Combining peer+random contexts resulted in a linear interpolation of the final losses of finetuning on either context. In other words, mixing random and peer context outperformed the final loss of finetuning exclusively on random context by a large margin. Here, we found evidence of additional predictive information in the peer context, which was not contained in the random context. However, mixing peer+user contexts did not result in a significantly lower final loss. This suggests a significant overlap in predictive information between peer and user context, upholding our claim that user context is strictly better than peer context.
4.4 User Tweets are Hard to Predict
In Figure 9 we present the distribution of the average model uncertainty on the tweets of our subjects. To gain an intuition on how hard it is to predict our subjects’ tweets relative to other accounts, we included three popular financial news accounts as comparison: YahooFinance, CNBC, and Stocktwits. Intuitively, predicting financial news and the stock market is hard Johnson et al. (2003), which should make them less predictable. However, Figure 9 reveals that most subjects in our pool are actually harder to predict than those news accounts. Of course, while the content of those tweets may be hard to predict, their style and vocabulary may not.
Finally, the absolute information content of users’ tweets is between 1.5-2 bits per character, depending on which model we use (see ’no context’ bars in Figure 2(c)). As a comparison, Shannon’s upper bound on the cross-entropy of the English language is 1.3 bits per character Shannon (1951), and recent estimates using neural language models say it is as low as 1.12 bits Takahashi and Tanaka-Ishii (2018). These results suggest that individual pieces of our online expression taken out of context are far from predictable using today’s LLMs. Provided with additional user context, only our largest model (Llama-2 70b) achieves comparable entropy, with an average of 1.1193 bits per character. While humans converge to 1.3 bits after only seeing $\sim$32 characters Moradi et al. (1998), language models typically need much longer context (see example from Appendix B.9).
Figure 9: Distribution of the average NLL of our subjects (with no additional context). As comparison we include the average NLL of three financial news accounts: YahooFinance, CNBC and Stocktwits. Model: GPT-2-XL.
@@ 5 Limitations
While negative log-likelihood is the most popular measure of model uncertainty, there are certain limitations to using this metric. One is that it is calculated exclusively on tokens contained in tweet $T$, and does not take into account any improvements on semantically equivalent text Kuhn et al. (2023). Another related issue is calibration Xiao et al. (2022). We also found the models to be highly sensitive to slight changes in prompting e.g. to how tweets were separated (App. A.3). Model uncertainty could also largely vary on the first token (App. B.6). Another important aspect is the length of input; NLL is commonly lower for longer pieces of text (App. B.3). While these points affect absolute numbers in our analysis, it does not affect our main statements about the relative comparisons between different models and contexts.
@@
Furthermore, we cannot rule out the possibility that some tweets in our corpus were part of the training dataset of the LLMs that we used. However, we believe the risk and severity of data contamination to be limited; we provide more detail as to why in Appendix A.2.
External validity
Extending our analysis to different sub-communities and languages would be an interesting avenue for future work. Similarly, we welcome additional work into the external validity of our results. It is unclear to what extent our findings would translate to other social media platforms.
@@ 6 Discussion
We presented the results of an investigation using state-of-the-art large language models into the predictability of online speech by analyzing posts on X (Twitter). As the basis of our study, we collected posts from more than five thousand users’ timelines and their peers. We used a total of 6.25M tweets for our main experiments, plus an additional 10M for tweet-tuning.
We found that it is surprisingly difficult to predict online speech. The majority of our users’ tweets are less predictable than tweets about financial news. Even with additional context, prediction performs relatively poorly. Only our largest model with additional user context is able to approximate the estimated entropy rate of the English language. Further analysis reveals that the source of greatest uncertainty for base models are hashtags and @-mentions. In other words, what improves prediction the most with additional context are basic signals: specifically, guessing the correct hashtags and @-mentions. All in all, our findings suggest that despite the impressive capabilities of large language models in other areas, state-of-the-art models predict online speech rather poorly.
@@
Our findings may also inform work on privacy and other ethical issues on social networks. For example, threats such as shadow profiling or impersonation on a global scale may not be as acute as some feared with state-of-the-art language models. However, these risks may still exist on a case-by-case basis (we demonstrate individual variability in Appendix B.2). Moreover, there may be other conceivable harms that don’t map cleanly to questions of predictability.
@@ A.1.1 Sampling stage
This collection phase resulted in a pool of 10M tweets, with 5M unique authors. We sampled our subjects from this pool of authors ($N=5102$ sample). Strictly speaking, our sample will be biased towards users that have 1) been more active during our initial collection period and 2) are more active users in general. This was the closest we could get to a random sample with the offered API endpoints. This practice follows Bagrow et al. (2019)’s method of randomly sampling users.
Twitter had roughly $\sim$500M active monthly users in 2023 Yaccarino (2023) – some of which may have been bot accounts Lee et al. (2011). To address this issue, we decided to filter out accounts that scored high (above 0.5 on a scale of 0 to 1) on the Bot-O-Meter bot detection API Sayyadiharikandeh et al. (2020). Bot-O-Meter uses classifies users on a scale from 0-1 based on 200 of their tweets. See Figure 10 for a distribution of these scores. We dropped users that had a score higher than 0.5. Additionally, we dropped users that had a high retweet ratio (more than $80\%$ of their tweets consisted of retweets). This is an additional measure to prevent bot accounts in our subject pool (bots are known for frequently retweeting content Yang et al. (2020); Gilani et al. (2017)) as well as a practical consideration since we only wanted to include non-retweets in our analysis.
@@ A.1.2 Timeline collection stage
We used the Twitter API’s Timeline endpoint to query the subjects’ most recent tweets. To get tweets from around the same timeframe, we choose an end_time (which was the start of our sampling stage, 20. January). Tweets made after end_time were not included. Per subject, we collected a total of 500 tweets, half of which was reserved for evaluation, the other half served as user context. Some user tweets date back as early as 2011, however the vast majority (95%) were authored after 2022.
From these tweets, we identified the top-15 most mentioned users (the user’s peers) and collected 50 tweets from their timelines as well. These, as well as the user context tweets were collected such that they were authored before the oldest evaluation tweet ($t^{*}$ in Figure 12). From these, we selected the 250 most recent tweets. A simpler approach would have been to select exactly $\frac{250}{15}\approx 17$ tweets per peer. However, not all peers had this many tweets on their timeline before $t^{*}$, hence our choice for collecting more tweets per peer than necessary.